### v1

# A Proxy-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

Abstract
Poisoning efficiency is a crucial factor in poisoning-based backdoor attacks. Attackers prefer to use as few poisoned samples as possible to achieve the same level of attack strength, in order to remain undetected. Efficient triggers have significantly improved poisoning efficiency, but there is still room for improvement. Recently, selecting efficient samples has shown promise, but it requires a proxy backdoor injection task to find an efficient poisoned sample set, which can lead to performance degradation if the proxy attack settings are different from the actual settings used by the victims. In this paper, we propose a novel Proxy-Free Strategy (PFS) that selects efficient poisoned samples based on individual similarity and set diversity, effectively addressing this issue. We evaluate the proposed strategy on several datasets, triggers, poisoning ratios, architectures, and training hyperparameters. Our experimental results demonstrate that PFS achieves higher backdoor attack strength while 500 faster than previous proxy-based selection approaches.
1 Introduction
The abundance of training data plays a vital role in the success of Deep Neural Networks (DNNs). For instance, GPT-3 [2], a deep learning model with 175 billion parameters, owes its effectiveness in various language processing tasks to its pretraining on a massive corpus of 45 TB of text data. With the increasing demand for data, users and businesses are turning to third-party sources or online collection as easier means of obtaining it. However, recent studies [4, 12, 3, 11] have demonstrated that such practices can be maliciously exploited by attackers to contaminate training data, thereby negatively impacting the functionality and reliability of trained models.
During the training phase, one of the significant threats is called backdoor attacks [4, 12, 27]. This attack involves injecting a covert backdoor into the DNN by introducing a small number of poisoned samples into the benign training set. Although the model appears normal on benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the attacker’s objective. As research continues to reveal this threat in various tasks such as speaker verification [46] and malware detection [20], attention to backdoor attacks is increasing.
Poisoning efficiency is a crucial indicator for attackers attempting to launch data poisoning attacks on DNNs [47, 22, 39, 40]. The aim is to contaminate the training dataset with as few poisoned samples as possible while achieving the desired outcome. A perfect scenario is if a backdoor can be successfully implanted by poisoning only one sample in the dataset, significantly enhancing the attack’s stealthiness. Currently, methods to improve poisoning efficiency fall into two categories: designing efficient triggers [47, 48, 44] and selecting efficient samples [39]. The former aims to find trigger patterns that are easier for models to learn, while the latter focuses on selecting and representative poisoned samples. This paper is concerned with the latter approach.
A recent study [39] on sample selection has yielded significant results, demonstrating that the same attack success rate can be achieved with only to of the poisoned samples compared to random selection strategies. However, we have found that this method requires a proxy backdoor injection task to evaluate the contribution of each poisoned sample, resulting in a constructed poisoned set closely related to the attacker’s specific proxy attack. This presents a challenge, as the effectiveness of the selected set will be damaged when the attacker’s proxy attack settings differ from the actual settings of the victims, which is a common scenario in practice. Moreover, conducting the proxy task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
In this paper, we address the above issue and present a Proxy-Free Strategy (PFS) that acquires higher backdoor attack strength while 500 faster than previous selection method. Our main contributions are as follows:
We conduct an analysis of the forensic features of efficient poisoned samples in backdoor attacks, with a specific focus on two key factors: the similarity between clean and corresponding poisoned samples, and the diversity of the poisoned sample set.
Based on this analysis, we propose a novel selection method called the Proxy-Free Strategy (PFS), which does not require a proxy attack task, thereby addressing the issue of inconsistent settings and significantly reducing the time consumption.
we evaluate PFS on three datasets: CIFAR-10, CIFAR-100, and Tiny-ImageNet, and our results consistently demonstrate its effectiveness. Notably, PFS achieves comparable or even higher attack success rates while 500 faster than previous proxy-based selection approaches, making it more scalable for larger datasets or complex models.
2 Related Works
2.1 Backdoor Attacks
Backdoor attacks aim to embed hidden Trojans into DNNs, enabling the attacked models to perform well on clean samples while misbehaving on samples with specifically designed triggers. These attacks can occur at any stage of DNN development, including code [1, 33], outsourcing, pretrained [36, 10, 19], data collection [12, 4, 35, 26], collaborative learning [42, 31], and post-deployment [32, 17]. Among these, poisoning-based backdoor attacks, which involve injecting backdoor triggers by modifying the training data, are the most straightforward and widely adopted. Recent studies have focused on improving poisoning efficiency, which can be divided into two categories.
Designing Efficient Trigger. Trigger design is a crucial aspect of backdoor attacks. Chen et al. [4] proposed a blended strategy that blends clean samples with the trigger to evade human detection. Later, studies aimed to create more efficient and invisible triggers by adopting natural patterns such as warping [30], rotation [38], style transfer [5], frequency [8, 45], and reflection [28]. In some studies, researchers have taken inspiration from Universal Adversarial Perturbations (UAPs) [29] and optimized an UAP on a pre-trained clean model to create a trigger, which is an effective and widely used approach [48, 21, 6]. Li et al. [23] proposed a different approach where they use GANs models to acquire a sample-specific trigger, which is a departure from previous works that adopt universal triggers.
Selecting Efficient Poisoning Samples. Efficient sample selection for poisoning attacks is a crucial but under-explored area, and is separate from the trigger design. Xia et al. [39] were among the first to explore the contribution of different data to backdoor injection, discovering that not all poisoned samples contribute equally and that appropriate sample selection can significantly improve the efficiency of data in backdoor attacks. However, their study [39] relies on proxy attacks to identify the efficient subset, which suffers from performance degradation when gaps exist (e.g., in Transform (F) and Hyperparameters (H) as shown in Figure 1 between the proxy poisoning attack and the actual poisoning process used by victims. Therefore, a sample selection strategy that is not based on proxy attacks is urgently needed, and is the main contribution of our paper.
2.2 Backdoor Defenses
Backdoor attacks pose a significant threat to the security of DNN models, and several defense methods have been proposed to mitigate these threats. Existing backdoor defenses can be broadly classified into two categories: backdoor detection and backdoor erasing. Backdoor detection aims to detect poisoning in input data or DNN models used [9, 14, 7, 41]. While these methods have demonstrated promising detection precision, they cannot completely mitigate the potential threats of backdoor models. Backdoor erasing aims to purify the threatened models by erasing the backdoor during or post training. During training, clean models can be trained on poisoned data, as proposed by Li et al. [24]. Post-training, models can be fine-tuned, distilled, or pruned to erase the backdoor from the attacked model [25, 37]. These defense methods offer varying levels of effectiveness, and the backdoor defense problem remains an active area of research.
3 Degradation of Proxy-based Selection
This section aims to highlight the limitations of sample selection methods that rely on proxy attacks, motivating the development of a proxy-free approach for selecting poisoned samples in backdoor attacks. To provide context, we will first review the general pipeline of backdoor attacks and the use of proxy attack-based sample selection methods
Poisoning-based Backdoor Attacks. Poisoning-based backdoor attacks involve injecting a backdoor into a subset of a clean training set , resulting in a corresponding poisoned set . Here, is a pre-defined poison generator that adds the trigger into a clean sample , while and represent the true label and attack-target label of the clean sample and the poisoning sample , respectively. After poisoning, the victim will train a deep model:
In the above equation, represents the DNN model, and is the corresponding loss function. The poisoned rate can be calculated as . The goal of sample selection methods is to enhance the efficiency of backdoor attacks by selecting an optimal poisoning set . This approach is orthogonal to the mainstream method of improving poisoning efficiency and has not been thoroughly explored yet.
Sample Selection Based on Proxy Attack. Recent studies [39, 43] have highlighted the varying contributions of different samples to backdoor injection, with certain subsets of efficient data exhibiting higher attack success rates. For instance, [39] suggests that the efficiency of a poisoned samples is related to its forgetting events during backdoor injection. To identify the optimal subset, they formulate sample selection as an optimization problem and propose a Filtering-and-Updating Strategy (FUS). Although FUS achieves remarkable performance, it utilizes a proxy attack processing to identify the efficient subset.
As illustrated in Figure 1, poisoning-based backdoor attacks consist of two phases: The attacker combines data (D) and trigger (T) to generate poisoned data (D+T) and releases it to victims, who then use it to train infected DNN models. During training, various data transforms/augmentations (F) and hyperparameters (H)***In this case, H includes any training settings such as the model architecture, the optimizer, and the training hyperparameters. are applied. Typically, the attacker has no access to any information about the victim phase. However, proxy-based sample selection methods rely on the complete attack process to select samples. We argue that all components involved in backdoor attacks, including D, T, F, and H, can affect the efficiency of poisoned samples. When discrepancies arise (e.g., between the transforms and hyperparameters used by the attacker and those used by victims), the effectiveness of the selected samples may be compromised. Therefore, we will first investigate the impact of such discrepancies on proxy-based sample selection methods.
3.1 Experimental Settings
To investigate the proposed question in this section, we conduct experiments using the VGG-16 model [34] as the victim model on the CIFAR-10 dataset [18]. We use the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, and apply two common data transformations: random crop and random horizontal flip. The training is performed for 70 epochs with a batch size of 256. Poisoned samples are generated using a blending strategy [4] in which the clean samples are blended with a trigger image using the equation , where is set to 0.15. We set the attack target to category 0. We adopt the FUS method with a poisoned rate (the size of the poisoning set is ) and the number of selection iterations following [39]. In accordance with previous literature [22], the attack success rate is defined as the probability of classifying a poisoned test data to the target label. Unless stated otherwise, all experiments in this section follow the aforementioned settings.
3.2 Empirical Studies
Different Data Transforms (F) will decrease the effectiveness of proxy-based sample selection method. We investigated how different data transforms (F) affect the effectiveness of proxy-based sample selection methods. To conduct our analysis, we used the experimental settings described previously and followed them exactly. Specifically, we set the data transform for the proxy poisoning attack in the sample selection phase (FUS) to ”RandomCrop” and ”RandomHorizontalFlip”, and we used ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter” for the data transform during the actual poisoning attack in the victim phase. Our results, depicted in Figure 2, indicate that different data transforms (F) decrease the effectiveness of proxy-based sample selection methods.
Different Hyperparameter (H) will decrease the effectiveness of proxy-based sample selection method. Furthermore, we explore the impact of different hyperparameter to efficient data searched by FUS [39]. Following the same experimental settings as described above, we conduct experiments using the hyperparameter ”V-16, SGD, 0.01” to search for efficient sets of backdoor attacks and evaluate the attack success rate on several other hyperparameters, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”†††V-13, V-16, R-18, and P-18 denote VGG13, VGG-16, ResNet-18, and PreActResNet-18, respectively. . Our results, shown in Figure 3, demonstrate that different hyperparameters (H) can also significantly decrease the effectiveness of the proxy-based sample selection method.
Overall, our results suggest that the efficacy of proxy-based sample selection methods is sensitive to attacker-agnostic data transforms and hyperparameters used during the victim phase. This highlights the importance of using proxy-free search methods to achieve efficient sample selection in backdoor attacks.
Input:
Clean training set ; Size of clean training set ; Backdoor trigger ; Attack target ; Poisoned rate ; Diversity rate ; Pre-trained feature extractor ; Trigger function
Output:
Build the poisoned set ;
4 Methodology
To improve poisoning efficiency, it is important to develop proxy-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks. Before introducing our method, we will first discuss some key observations.
4.1 Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
We contend that the primary reason for the efficiency of backdoor injection lies in the similarity between clean and corresponding poisoned samples. Typically, the labels of poisoned samples are different from those of the original clean samples, meaning that clean and corresponding poisoned samples share similar features but have completely different labels. In this case, samples with high similarity can be viewed as challenging samples in the poisoning task, making them more efficient than low-similarity samples for backdoor attacks. To validate this hypothesis, we conducted empirical studies on the CIFAR-10 dataset using the same experimental settings as in Sec. 3.1. The left part of Figure 4 depicts the similarity distribution of 500 randomly selected samples and 500 FUS-searched samples [39], which demonstrates that FUS-searched efficient data has higher similarity than random samples. We further examined the contribution of similarity to the data efficiency of backdoor attacks by selecting poisoned samples based on similarity. The right part of Figure 4 shows the attack success rate when different poisoned samples were used. As expected, poisoning with high-similarity samples (Top-500 similarity sampling) resulted in a significantly higher attack success rate than poisoning with low-similarity samples (Bottom-500 similarity sampling). Furthermore, we compared random sampling with similarity-selected sampling. However, both poisoning with high-similarity samples (Top-500 similarity sampling) and low-similarity samples (Bottom-500 similarity sampling) yielded lower attack success rates than poisoning with random samples (Random sampling). This suggests that similarity can only be used to filter out inefficient samples and has no positive correlation with data efficiency in poisoned attacks. We argue that the degradation of Top-500 similarity samples can be attributed to the overly limited diversity of the poisoned samples. The middle part of Figure 4 shows the t-SNE of the 50K whole dataset and the 500 selected samples, indicating that Top-500 similarity samples have limited diversity. Selecting 500 samples randomly from the 5K samples with the highest similarity (Top-5K similarity random 500 sampling) can effectively mitigate this limitation, as shown in the right part of Figure 4. Among them, Top-5K similarity random 500 sampling outperforms random sampling and achieves competitive results with only a slight cost compared to FUS [39] methods. In summary, the similarity between clean and corresponding poisoned samples and the diversity of poisoned sample set are two key factors in the efficient data of backdoor attacks.
4.2 Threat Model
Attacker’s Capacities. Following the basic requirement of poisoning-based backdoor attacks [22, 39], we assume that attackers are permitted to poison a portion of the training data, but are not privy to any other training components in the victim phase, such as the training loss, schedule, or model structure. These assumptions align with potential real-world scenarios in which such threats can occur.
Attacker’s Goals. Our aim is to enhance the efficiency of poisoning attacks without incurring additional costs, by selecting an appropriate poisoning set from the dataset . This approach is distinct from the current methods that enhance efficiency by designing triggers, and hence, our method can be easily incorporated into other attack technologies at minimal cost.
4.3 Proxy-Free Selection Strategy
Motivated by the observation of forensic features that contribute to efficient data in poisoning-based backdoor attacks, as demonstrated in Section 4.1, we propose a simple yet effective sample selection strategy (PFS) for enhancing poisoning efficiency. Unlike previous studies that rely on a proxy backdoor attack for sample selection, our method is based on the similarity between clean and corresponding poisoned samples, which is only dependent on the Data (D) and Trigger (T), and independent of Data Transform (F) and Hyperparameters (H) (proxy-free). Specifically, we utilize a pre-trained feature extractor to compute the cosine similarity (cos()) in the feature space between clean and corresponding poisoned samples. We then select the top most similar samples and randomly sample from this set as the poisoning set, where is a hyper-parameter that controls the diversity of the poisoned samples. The algorithmic procedure of our method is presented in Algorithm 1.
Moreover, we demonstrate that our method can filter out most of the worthless samples, which also enhances the FUS search [39]. Therefore, we propose a combined approach (FUS+PFS) that utilizes the FUS search on the coarse poisoned set containing samples to further enhance the poisoning efficiency.
5 Experiments
We evaluate the effectiveness of PFS on datasets: CIFAR -10 (Sec. 5.2), CIFAR-100 (Sec. 5.3), and Tiny-ImageNet (Sec. A.4 of the Supplementary Materials). We also provide experimental settings in Sec. 5.1, time-consuming analysis in Sec. 5.4, and ablation study in Sec. 5.5.
5.1 Experimental Settings
We conduct experiments with three trigger types (BadNets [13], Blended [4], Optimized [48]), six data transforms (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three DNN architectures (VGG16 [34], ResNet18 [15], PreAct-ResNet18 [16]). We also use four optimizers: SGD-0.01‡‡‡SGD-0.01 means adopting the SGD optimizer with an initial learning rate of 0.01,, SGD-0.02, Adam-0.001, Adam-0.002. The adopted feature extractor is the ResNet-18 that has been pre-trained on the current dataset. Additionally, in Section A.5 of the Supplementary Materials, we present experiments with different pre-trained feature extractors on the CIFAR-10 dataset. Our results show that even using a common feature extractor pre-trained on the ImageNet dataset yields significant improvements compared to the baseline.
We compare our method to the commonly used random selection (Random) [13, 4] and FUS [39]. In the FUS search process, we use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. For our method, we use pre-trained ResNet models for similarity measurement and set the diversity hyperparameter to 10 for all three datasets. Other hyperparameters are kept the same as in Sec. 3.1.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.950 | 0.984 | 0.978 | 0.988 | 0.964 | 0.992 | 0.988 | 0.994 | 0.961 | 0.991 | 0.985 | 0.994 |
| RandomCrop | 0.906 | 0.964 | 0.954 | 0.972 | 0.918 | 0.976 | 0.966 | 0.976 | 0.927 | 0.977 | 0.968 | 0.981 | |
| RandomHorizontalFlip | 0.932 | 0.981 | 0.971 | 0.986 | 0.964 | 0.989 | 0.983 | 0.992 | 0.954 | 0.987 | 0.981 | 0.990 | |
| RandomRotation | 0.879 | 0.952 | 0.939 | 0.953 | 0.909 | 0.972 | 0.955 | 0.976 | 0.918 | 0.974 | 0.961 | 0.977 | |
| ColorJitter | 0.952 | 0.982 | 0.975 | 0.987 | 0.967 | 0.993 | 0.988 | 0.995 | 0.964 | 0.992 | 0.985 | 0.994 | |
| RandomCrop+ RandomHorizontalFlip | 0.916 | 0.966 | 0.956 | 0.972 | 0.933 | 0.980 | 0.970 | 0.983 | 0.931 | 0.980 | 0.970 | 0.983 | |
| Avg | 0.923 | 0.972 | 0.962 | 0.976 | 0.943 | 0.984 | 0.975 | 0.986 | 0.943 | 0.984 | 0.975 | 0.987 | |
| Blended | None | 0.795 | 0.856 | 0.861 | 0.889 | 0.814 | 0.923 | 0.885 | 0.930 | 0.867 | 0.949 | 0.919 | 0.957 |
| RandomCrop | 0.792 | 0.830 | 0.865 | 0.874 | 0.818 | 0.888 | 0.885 | 0.915 | 0.821 | 0.903 | 0.896 | 0.921 | |
| RandomHorizontalFlip | 0.827 | 0.899 | 0.916 | 0.930 | 0.867 | 0.948 | 0.925 | 0.959 | 0.864 | 0.954 | 0.923 | 0.962 | |
| RandomRotation | 0.791 | 0.821 | 0.855 | 0.864 | 0.833 | 0.897 | 0.890 | 0.926 | 0.838 | 0.918 | 0.902 | 0.933 | |
| ColorJitter | 0.758 | 0.846 | 0.826 | 0.861 | 0.758 | 0.846 | 0.826 | 0.921 | 0.758 | 0.846 | 0.826 | 0.950 | |
| RandomCrop+ RandomHorizontalFlip | 0.827 | 0.883 | 0.911 | 0.933 | 0.855 | 0.931 | 0.932 | 0.958 | 0.864 | 0.926 | 0.924 | 0.956 | |
| Avg | 0.798 | 0.856 | 0.872 | 0.892 | 0.824 | 0.906 | 0.891 | 0.935 | 0.835 | 0.916 | 0.898 | 0.947 | |
| Optimized | None | 0.844 | 0.973 | 0.910 | 0.988 | 0.840 | 0.952 | 0.897 | 0.968 | 0.902 | 0.982 | 0.940 | 0.993 |
| RandomCrop | 0.952 | 0.998 | 0.990 | 1.000 | 0.966 | 0.999 | 0.994 | 1.000 | 0.970 | 0.999 | 0.993 | 1.000 | |
| RandomHorizontalFlip | 0.922 | 0.990 | 0.975 | 0.998 | 0.885 | 0.977 | 0.945 | 0.993 | 0.932 | 0.994 | 0.981 | 0.998 | |
| RandomRotation | 0.844 | 0.893 | 0.873 | 0.874 | 0.828 | 0.930 | 0.879 | 0.914 | 0.863 | 0.958 | 0.918 | 0.956 | |
| ColorJitter | 0.875 | 0.962 | 0.933 | 0.991 | 0.861 | 0.940 | 0.918 | 0.977 | 0.911 | 0.985 | 0.969 | 0.995 | |
| RandomCrop+ RandomHorizontalFlip | 0.979 | 0.999 | 0.996 | 1.000 | 0.979 | 0.999 | 0.997 | 1.000 | 0.982 | 0.999 | 0.997 | 1.000 | |
| Avg | 0.903 | 0.969 | 0.946 | 0.975 | 0.893 | 0.966 | 0.938 | 0.975 | 0.927 | 0.986 | 0.966 | 0.990 | |
5.2 Experiments on CIFAR-10 Dataset
Results on different data transforms and architectures. Table 1 presents the attack success rate for BadNets, Blended, and Optimized triggers using different data transforms and architectures. We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. Our PFS method significantly improves the efficiency of poisoned data for backdoor attacks. In all 54 settings, the attack success rate of the poisoned samples selected by our method is higher than that of randomly selected samples, at the same poisoning ratios. On average, our method outperforms random selection by 0.044, 0.074, and 0.066 for BadNets, Blended, and Optimized triggers, respectively. In general, our method performs better than FUS, achieving stronger attack strength in 48 out of 54 settings. It is worth noting that the FUS search requires hundreds or even thousands of times more time than our method. Combining FUS with our method yields the best results in 51 out of 54 settings, as our method filters out many inefficient samples and narrows the FUS search space.
Results on different optimizer and architectures. Table 2 shows the attack success rate for the BadNets, Blended, and Optimized triggers with different optimizers and architectures. We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. The results demonstrate the effectiveness of our PFS method. First, for all 36 settings, the attack success rate achieved by our method is higher than that achieved by random selection, with average improvements of 0.036, 0.0437, and 0.017 for the BadNets, Blended, and Optimized triggers, respectively. Second, our method outperforms FUS in 25 out of 36 cases. When combining our method with FUS, we obtain the best results in 33 out of 36 settings.
In the Supplementary Materials, we have included additional experiments to further evaluate the effectiveness and robustness of our proposed strategy. Specifically, in Sec. A.1 and A.2, we present the attack success rate (ASR) and benign accuracy (BA) under poisoned rates of and on the CIFAR-10 dataset. The results demonstrate that our method remains robust to different poisoned rates and target labels while preserving benign accuracy. Additionally, in Sec. A.3, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset. Overall, our findings confirm the effectiveness of our proposed method and support our hypothesis that the efficiency of backdoor attacks is determined by sample similarity and diversity.
| Trigger | Optimizer | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | SGD-0.01 | 0.916 | 0.966 | 0.956 | 0.972 | 0.933 | 0.980 | 0.970 | 0.983 | 0.931 | 0.980 | 0.970 | 0.983 |
| SGD-0.02 | 0.918 | 0.971 | 0.962 | 0.976 | 0.938 | 0.983 | 0.975 | 0.986 | 0.943 | 0.984 | 0.976 | 0.987 | |
| Adam-0.0005 | 0.920 | 0.934 | 0.963 | 0.971 | 0.945 | 0.960 | 0.975 | 0.985 | 0.944 | 0.960 | 0.976 | 0.984 | |
| Adam-0.001 | 0.928 | 0.971 | 0.967 | 0.957 | 0.950 | 0.985 | 0.977 | 0.986 | 0.954 | 0.982 | 0.977 | 0.985 | |
| Avg | 0.921 | 0.961 | 0.962 | 0.969 | 0.942 | 0.977 | 0.974 | 0.985 | 0.943 | 0.977 | 0.975 | 0.985 | |
| Blended | SGD-0.01 | 0.827 | 0.883 | 0.911 | 0.933 | 0.855 | 0.931 | 0.932 | 0.958 | 0.864 | 0.926 | 0.924 | 0.956 |
| SGD-0.02 | 0.836 | 0.887 | 0.934 | 0.935 | 0.881 | 0.939 | 0.950 | 0.968 | 0.881 | 0.947 | 0.948 | 0.967 | |
| Adam-0.0005 | 0.873 | 0.890 | 0.924 | 0.908 | 0.897 | 0.952 | 0.953 | 0.974 | 0.893 | 0.952 | 0.940 | 0.966 | |
| Adam-0.001 | 0.886 | 0.910 | 0.937 | 0.909 | 0.900 | 0.953 | 0.951 | 0.970 | 0.896 | 0.949 | 0.943 | 0.962 | |
| Avg | 0.881 | 0.891 | 0.927 | 0.921 | 0.883 | 0.944 | 0.947 | 0.968 | 0.884 | 0.944 | 0.939 | 0.963 | |
| Optimized | SGD-0.01 | 0.979 | 0.999 | 0.996 | 1.000 | 0.979 | 0.999 | 0.997 | 1.000 | 0.982 | 0.999 | 0.997 | 1.000 |
| SGD-0.02 | 0.979 | 0.999 | 0.998 | 1.000 | 0.982 | 0.999 | 0.997 | 1.000 | 0.984 | 0.999 | 0.997 | 1.000 | |
| Adam-0.0005 | 0.975 | 0.997 | 0.996 | 1.000 | 0.986 | 0.998 | 0.996 | 1.000 | 0.982 | 0.998 | 0.996 | 1.000 | |
| Adam-0.001 | 0.980 | 0.999 | 0.997 | 1.000 | 0.985 | 0.999 | 0.997 | 1.000 | 0.982 | 0.998 | 0.994 | 1.000 | |
| Avg | 0.978 | 0.999 | 0.997 | 1.000 | 0.983 | 0.999 | 0.997 | 1.000 | 0.983 | 0.997 | 0.996 | 1.000 | |
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.933 | 0.970 | 0.965 | 0.971 | 0.950 | 0.978 | 0.973 | 0.979 | 0.949 | 0.977 | 0.970 | 0.977 |
| RandomCrop | 0.882 | 0.947 | 0.930 | 0.945 | 0.894 | 0.951 | 0.940 | 0.952 | 0.888 | 0.951 | 0.940 | 0.951 | |
| RandomHorizontalFlip | 0.920 | 0.964 | 0.958 | 0.965 | 0.936 | 0.970 | 0.965 | 0.971 | 0.929 | 0.968 | 0.962 | 0.968 | |
| RandomRotation | 0.868 | 0.923 | 0.912 | 0.924 | 0.866 | 0.939 | 0.925 | 0.938 | 0.875 | 0.935 | 0.926 | 0.938 | |
| ColorJitter | 0.934 | 0.970 | 0.965 | 0.971 | 0.952 | 0.978 | 0.973 | 0.979 | 0.946 | 0.977 | 0.970 | 0.978 | |
| RandomCrop+ RandomHorizontalFlip | 0.879 | 0.939 | 0.929 | 0.939 | 0.888 | 0.947 | 0.942 | 0.950 | 0.894 | 0.949 | 0.941 | 0.951 | |
| Avg | 0.903 | 0.952 | 0.943 | 0.953 | 0.914 | 0.961 | 0.953 | 0.962 | 0.914 | 0.960 | 0.952 | 0.961 | |
| Blended | None | 0.866 | 0.923 | 0.920 | 0.941 | 0.884 | 0.963 | 0.946 | 0.974 | 0.894 | 0.975 | 0.950 | 0.982 |
| RandomCrop | 0.863 | 0.874 | 0.891 | 0.880 | 0.826 | 0.916 | 0.906 | 0.938 | 0.829 | 0.931 | 0.912 | 0.942 | |
| RandomHorizontalFlip | 0.891 | 0.945 | 0.935 | 0.960 | 0.902 | 0.968 | 0.951 | 0.977 | 0.912 | 0.974 | 0.958 | 0.982 | |
| RandomRotation | 0.818 | 0.862 | 0.901 | 0.865 | 0.841 | 0.916 | 0.916 | 0.936 | 0.848 | 0.926 | 0.921 | 0.944 | |
| ColorJitter | 0.863 | 0.920 | 0.917 | 0.931 | 0.904 | 0.972 | 0.956 | 0.977 | 0.913 | 0.979 | 0.959 | 0.983 | |
| RandomCrop+ RandomHorizontalFlip | 0.811 | 0.865 | 0.905 | 0.872 | 0.826 | 0.918 | 0.920 | 0.942 | 0.850 | 0.931 | 0.924 | 0.947 | |
| Avg | 0.852 | 0.898 | 0.912 | 0.908 | 0.864 | 0.942 | 0.933 | 0.957 | 0.874 | 0.953 | 0.937 | 0.963 | |
| Optimized | None | 0.979 | 0.999 | 0.995 | 0.999 | 0.884 | 0.972 | 0.931 | 0.982 | 0.895 | 0.983 | 0.962 | 0.987 |
| RandomCrop | 0.989 | 0.999 | 0.997 | 1.000 | 0.983 | 0.999 | 0.996 | 1.000 | 0.981 | 0.999 | 0.995 | 1.000 | |
| RandomHorizontalFlip | 0.985 | 0.999 | 0.997 | 1.000 | 0.938 | 0.998 | 0.980 | 0.983 | 0.938 | 0.997 | 0.984 | 0.998 | |
| RandomRotation | 0.810 | 0.831 | 0.867 | 0.785 | 0.827 | 0.909 | 0.882 | 0.882 | 0.835 | 0.908 | 0.883 | 0.900 | |
| ColorJitter | 0.984 | 0.997 | 0.993 | 1.000 | 0.967 | 0.989 | 0.979 | 0.985 | 0.956 | 0.993 | 0.990 | 0.994 | |
| RandomCrop+ RandomHorizontalFlip | 0.991 | 0.999 | 0.997 | 1.000 | 0.987 | 0.999 | 0.998 | 1.000 | 0.982 | 0.998 | 0.994 | 0.999 | |
| Avg | 0.956 | 0.971 | 0.974 | 0.964 | 0.931 | 0.978 | 0.961 | 0.972 | 0.931 | 0.980 | 0.968 | 0.980 | |
| PFS (w/o pre-training) | PFS (w pre-training) | FUS () |
| 17 | 562 | 9,150 |
5.3 Experiments on CIFAR-100 Dataset
The results for CIFAR-100 in Table 3 are similar to those for CIFAR-10. Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively. Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
5.4 Analysis of Time Consumption
Table 4 presents a comparison of the running times of our PFS with those of FUS. Our method demonstrates remarkable speed, being faster than FUS. Even with pre-training time included, our method is still approximately faster. Importantly, our method utilizes the same pre-trained model for different triggers, while FUS needs to conduct a complete search from the beginning for each trigger. This advantage makes our method a practical choice for situations where efficiency is crucial.
5.5 Ablation Study
We conduct an ablation study on the hyperparameter , which determines the diversity of the poisoned sample set. The results are shown in Figure 5, and the attack success rates demonstrate similar trends under different data transforms. Initially, as diversity gradually increases, the attack success rates also increase. However, too large of a value for can reduce the role of similarity and weaken the attack. These findings highlight the importance of balancing similarity and diversity for efficient poisoned samples, supporting our claim in Sec. 4.1.
6 Conclusion and Future Work
Conclusion. This paper presents empirical evidence and reveals two intriguing phenomena related to data efficiency in backdoor attacks. First, we demonstrate that sample selection methods based on proxy attacks suffer from performance degradation when gaps exist between proxy poisoning attacks and actual victim poisoning processes. Second, we show that the similarity between clean and corresponding poisoned samples, as well as the diversity of the poisoned sample set, are key features for efficient data in backdoor attacks. Based on these findings, we propose a simple yet effective proxy-free sample selection strategy that selects poisoned samples based on the similarity between clean and corresponding poisoned samples. Our method significantly improves the poisoning efficiency on benchmark datasets with almost no additional cost.
Future Work. While the proxy-free sampling strategy we proposed has demonstrated efficient data selection in backdoor attacks, we note that the similarity between clean and corresponding poisoned samples does not necessarily guarantee high efficiency, as similarity only filters out non-efficient samples. In the future, it would be desirable to identify a unified indicator that positively correlates with data efficiency in backdoor attacks.
Appendix A Supplementary Materials
This supplementary material is organized as follows. We first discuss the attack success rate and benign accuracy under poisoned rates with and on the CIFAR-10 dataset in Sec. A.1 and Sec. A.2, respectively. Additionally, we presents the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec. A.3. Besides, Sec. A.4 presents the experimental results on the Tiny-ImageNet dataset. At last, Sec. A.5 shows the experiments on different pre-trained feature extractor .
A.1 Attack Success Rate Under Additional Poisoned Rates on the CIFAR-10 Dataset
We evaluate the attack success rate (ASR) under different poisoned rates on the CIFAR-10 dataset. Tables 5 and 6 present the ASR with trigger BadNets and Blended at poisoned rates of and , respectively. Results under poisoned rates of and are similar to those at .
(i) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.030 and 0.055 on BadNets and Blended triggers, respectively. The proxy-free method also achieves better attack results than FUS in most cases (27/36). The combination of these two methods achieves superior results in all 36 settings.
(ii) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.023 and 0.04 on BadNets and Blended triggers, respectively. The proxy-free method also achieves better attack results than FUS in most cases (26/36). The combination of these two methods achieves superior results in all 36 settings.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.961 | 0.986 | 0.982 | 0.992 | 0.976 | 0.993 | 0.990 | 0.997 | 0.973 | 0.992 | 0.989 | 0.997 |
| RandomCrop | 0.924 | 0.970 | 0.966 | 0.983 | 0.944 | 0.980 | 0.974 | 0.988 | 0.946 | 0.979 | 0.977 | 0.989 | |
| RandomHorizontalFlip | 0.952 | 0.983 | 0.979 | 0.991 | 0.973 | 0.975 | 0.986 | 0.995 | 0.967 | 0.988 | 0.984 | 0.994 | |
| RandomRotation | 0.915 | 0.968 | 0.958 | 0.977 | 0.929 | 0.979 | 0.971 | 0.988 | 0.941 | 0.979 | 0.973 | 0.988 | |
| ColorJitter | 0.958 | 0.986 | 0.981 | 0.992 | 0.977 | 0.994 | 0.990 | 0.997 | 0.972 | 0.993 | 0.988 | 0.996 | |
| RandomCrop+ RandomHorizontalFlip | 0.932 | 0.970 | 0.965 | 0.981 | 0.949 | 0.980 | 0.978 | 0.989 | 0.952 | 0.981 | 0.978 | 0.989 | |
| Avg | 0.940 | 0.977 | 0.972 | 0.986 | 0.958 | 0.984 | 0.982 | 0.992 | 0.959 | 0.985 | 0.982 | 0.992 | |
| Blended | None | 0.866 | 0.921 | 0.918 | 0.953 | 0.883 | 0.954 | 0.939 | 0.971 | 0.909 | 0.970 | 0.954 | 0.982 |
| RandomCrop | 0.855 | 0.900 | 0.933 | 0.942 | 0.879 | 0.935 | 0.943 | 0.963 | 0.882 | 0.948 | 0.938 | 0.967 | |
| RandomHorizontalFlip | 0.889 | 0.937 | 0.952 | 0.970 | 0.906 | 0.966 | 0.960 | 0.983 | 0.927 | 0.975 | 0.961 | 0.987 | |
| RandomRotation | 0.859 | 0.909 | 0.929 | 0.937 | 0.894 | 0.938 | 0.947 | 0.967 | 0.901 | 0.955 | 0.953 | 0.972 | |
| ColorJitter | 0.851 | 0.915 | 0.912 | 0.938 | 0.872 | 0.950 | 0.929 | 0.969 | 0.913 | 0.966 | 0.945 | 0.976 | |
| RandomCrop+ RandomHorizontalFlip | 0.885 | 0.931 | 0.952 | 0.965 | 0.905 | 0.958 | 0.964 | 0.982 | 0.919 | 0.967 | 0.967 | 0.984 | |
| Avg | 0.868 | 0.919 | 0.933 | 0.951 | 0.890 | 0.950 | 0.947 | 0.973 | 0.909 | 0.964 | 0.953 | 0.978 | |
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.971 | 0.987 | 0.986 | 0.993 | 0.982 | 0.994 | 0.992 | 0.998 | 0.977 | 0.993 | 0.991 | 0.997 |
| RandomCrop | 0.940 | 0.973 | 0.970 | 0.986 | 0.954 | 0.980 | 0.979 | 0.991 | 0.946 | 0.982 | 0.979 | 0.992 | |
| RandomHorizontalFlip | 0.961 | 0.984 | 0.981 | 0.992 | 0.976 | 0.991 | 0.988 | 0.995 | 0.975 | 0.989 | 0.986 | 0.995 | |
| RandomRotation | 0.931 | 0.969 | 0.966 | 0.986 | 0.948 | 0.977 | 0.977 | 0.992 | 0.946 | 0.981 | 0.979 | 0.991 | |
| ColorJitter | 0.966 | 0.986 | 0.984 | 0.993 | 0.982 | 0.993 | 0.992 | 0.998 | 0.981 | 0.992 | 0.990 | 0.997 | |
| RandomCrop+ RandomHorizontalFlip | 0.938 | 0.974 | 0.972 | 0.986 | 0.957 | 0.982 | 0.980 | 0.992 | 0.959 | 0.984 | 0.981 | 0.991 | |
| Avg | 0.951 | 0.979 | 0.977 | 0.989 | 0.967 | 0.986 | 0.985 | 0.994 | 0.964 | 0.987 | 0.984 | 0.994 | |
| Blended | None | 0.914 | 0.943 | 0.952 | 0.952 | 0.916 | 0.969 | 0.955 | 0.981 | 0.937 | 0.978 | 0.967 | 0.988 |
| RandomCrop | 0.896 | 0.937 | 0.958 | 0.966 | 0.912 | 0.961 | 0.961 | 0.978 | 0.913 | 0.960 | 0.965 | 0.985 | |
| RandomHorizontalFlip | 0.923 | 0.955 | 0.972 | 0.983 | 0.931 | 0.976 | 0.973 | 0.990 | 0.947 | 0.982 | 0.976 | 0.991 | |
| RandomRotation | 0.900 | 0.936 | 0.952 | 0.963 | 0.915 | 0.958 | 0.958 | 0.973 | 0.922 | 0.965 | 0.965 | 0.982 | |
| ColorJitter | 0.904 | 0.940 | 0.951 | 0.963 | 0.918 | 0.965 | 0.949 | 0.979 | 0.938 | 0.973 | 0.964 | 0.986 | |
| RandomCrop+ RandomHorizontalFlip | 0.922 | 0.955 | 0.973 | 0.986 | 0.927 | 0.968 | 0.977 | 0.991 | 0.933 | 0.973 | 0.979 | 0.992 | |
| Avg | 0.910 | 0.944 | 0.960 | 0.969 | 0.920 | 0.966 | 0.962 | 0.969 | 0.932 | 0.972 | 0.969 | 0.987 | |
A.2 Benign Accuracy Under Additional Poisoned Rates on the CIFAR-10 Dataset
We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoned rates on the CIFAR-10 dataset. Table 7 and Table 8 illustrate the benign accuracy with trigger BadNets and Blended under poisoned rates of and , respectively. The results show the similar BA with four strategies, which demonstrates that our proposed strategy has no negative impact on benign accuracy.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.868 | 0.868 | 0.865 | 0.866 | 0.837 | 0.837 | 0.838 | 0.835 | 0.850 | 0.849 | 0.848 | 0.847 |
| RandomCrop | 0.903 | 0.905 | 0.905 | 0.903 | 0.914 | 0.912 | 0.914 | 0.911 | 0.914 | 0.914 | 0.914 | 0.914 | |
| RandomHorizontalFlip | 0.892 | 0.892 | 0.893 | 0.894 | 0.878 | 0.879 | 0.879 | 0.878 | 0.886 | 0.888 | 0.886 | 0.883 | |
| RandomRotation | 0.876 | 0.873 | 0.876 | 0.874 | 0.883 | 0.881 | 0.880 | 0.880 | 0.889 | 0.888 | 0.889 | 0.889 | |
| ColorJitter | 0.860 | 0.860 | 0.860 | 0.859 | 0.833 | 0.829 | 0.833 | 0.830 | 0.843 | 0.844 | 0.841 | 0.843 | |
| RandomCrop+ RandomHorizontalFlip | 0.920 | 0.919 | 0.920 | 0.921 | 0.927 | 0.926 | 0.928 | 0.925 | 0.927 | 0.928 | 0.929 | 0.927 | |
| Avg | 0.887 | 0.886 | 0.887 | 0.886 | 0.879 | 0.877 | 0.879 | 0.877 | 0.885 | 0.885 | 0.885 | 0.884 | |
| Blended | None | 0.865 | 0.866 | 0.865 | 0.866 | 0.836 | 0.832 | 0.832 | 0.833 | 0.846 | 0.846 | 0.849 | 0.847 |
| RandomCrop | 0.904 | 0.906 | 0.904 | 0.904 | 0.912 | 0.913 | 0.912 | 0.913 | 0.913 | 0.914 | 0.915 | 0.914 | |
| RandomHorizontalFlip | 0.892 | 0.892 | 0.889 | 0.891 | 0.876 | 0.878 | 0.875 | 0.875 | 0.882 | 0.884 | 0.885 | 0.887 | |
| RandomRotation | 0.873 | 0.875 | 0.875 | 0.873 | 0.880 | 0.880 | 0.881 | 0.880 | 0.887 | 0.888 | 0.889 | 0.889 | |
| ColorJitter | 0.858 | 0.859 | 0.860 | 0.858 | 0.829 | 0.832 | 0.824 | 0.828 | 0.839 | 0.841 | 0.843 | 0.842 | |
| RandomCrop+ RandomHorizontalFlip | 0.922 | 0.920 | 0.919 | 0.920 | 0.928 | 0.927 | 0.928 | 0.928 | 0.928 | 0.927 | 0.929 | 0.928 | |
| Avg | 0.886 | 0.886 | 0.885 | 0.885 | 0.877 | 0.877 | 0.875 | 0.876 | 0.883 | 0.883 | 0.885 | 0.885 | |
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.867 | 0.865 | 0.866 | 0.863 | 0.837 | 0.835 | 0.835 | 0.833 | 0.851 | 0.846 | 0.848 | 0.847 |
| RandomCrop | 0.904 | 0.905 | 0.905 | 0.905 | 0.913 | 0.914 | 0.912 | 0.912 | 0.916 | 0.915 | 0.913 | 0.913 | |
| RandomHorizontalFlip | 0.891 | 0.891 | 0.891 | 0.891 | 0.878 | 0.878 | 0.878 | 0.876 | 0.886 | 0.885 | 0.886 | 0.885 | |
| RandomRotation | 0.873 | 0.876 | 0.874 | 0.872 | 0.882 | 0.882 | 0.881 | 0.880 | 0.888 | 0.889 | 0.889 | 0.890 | |
| ColorJitter | 0.862 | 0.857 | 0.857 | 0.859 | 0.830 | 0.827 | 0.829 | 0.831 | 0.844 | 0.843 | 0.843 | 0.841 | |
| RandomCrop+ RandomHorizontalFlip | 0.920 | 0.920 | 0.919 | 0.919 | 0.928 | 0.926 | 0.926 | 0.925 | 0.928 | 0.930 | 0.928 | 0.927 | |
| Avg | 0.886 | 0.886 | 0.885 | 0.885 | 0.878 | 0.877 | 0.877 | 0.876 | 0.886 | 0.885 | 0.885 | 0.884 | |
| Blended | None | 0.865 | 0.866 | 0.865 | 0.866 | 0.836 | 0.832 | 0.832 | 0.833 | 0.846 | 0.846 | 0.849 | 0.847 |
| RandomCrop | 0.904 | 0.906 | 0.904 | 0.904 | 0.912 | 0.913 | 0.912 | 0.913 | 0.913 | 0.914 | 0.915 | 0.914 | |
| RandomHorizontalFlip | 0.892 | 0.892 | 0.889 | 0.891 | 0.876 | 0.878 | 0.875 | 0.875 | 0.882 | 0.884 | 0.885 | 0.887 | |
| RandomRotation | 0.873 | 0.875 | 0.875 | 0.873 | 0.880 | 0.880 | 0.881 | 0.880 | 0.887 | 0.888 | 0.889 | 0.889 | |
| ColorJitter | 0.858 | 0.859 | 0.860 | 0.858 | 0.829 | 0.832 | 0.824 | 0.828 | 0.839 | 0.841 | 0.843 | 0.842 | |
| RandomCrop+ RandomHorizontalFlip | 0.922 | 0.920 | 0.919 | 0.920 | 0.928 | 0.927 | 0.928 | 0.928 | 0.928 | 0.927 | 0.929 | 0.928 | |
| Avg | 0.886 | 0.886 | 0.885 | 0.885 | 0.877 | 0.877 | 0.875 | 0.876 | 0.883 | 0.883 | 0.885 | 0.885 | |
A.3 Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
This section presents the attack success rate (ASR) under additional target classes on the CIFAR-10 dataset. Table 9 illustrates the attack success rate with trigger BadNets and Blended under target category 3. Compared with random selection, our proxy-free approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers. The proxy-free method also achieves better attack results in most cases compared to FUS (22/24). The combination of these two achieves superior results in all 24 settings. These outcomes consistently confirm the conclusions we reached in the paper.
| Trigger | Data Transform | Model | |||||||
| VGG16 | ResNet18 | ||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.950 | 0.984 | 0.977 | 0.986 | 0.959 | 0.991 | 0.985 | 0.994 |
| RandomCrop | 0.909 | 0.960 | 0.954 | 0.968 | 0.925 | 0.976 | 0.966 | 0.979 | |
| RandomHorizontalFlip | 0.934 | 0.980 | 0.972 | 0.983 | 0.959 | 0.987 | 0.981 | 0.990 | |
| RandomRotation | 0.890 | 0.951 | 0.942 | 0.951 | 0.912 | 0.968 | 0.955 | 0.972 | |
| ColorJitter | 0.945 | 0.983 | 0.975 | 0.986 | 0.966 | 0.991 | 0.985 | 0.993 | |
| RandomCrop+ RandomHorizontalFlip | 0.905 | 0.964 | 0.957 | 0.968 | 0.925 | 0.978 | 0.968 | 0.980 | |
| Avg | 0.922 | 0.970 | 0.963 | 0.974 | 0.941 | 0.982 | 0.973 | 0.985 | |
| Blended | None | 0.774 | 0.867 | 0.851 | 0.899 | 0.806 | 0.915 | 0.850 | 0.935 |
| RandomCrop | 0.780 | 0.813 | 0.822 | 0.869 | 0.791 | 0.902 | 0.827 | 0.918 | |
| RandomHorizontalFlip | 0.836 | 0.899 | 0.887 | 0.934 | 0.845 | 0.949 | 0.888 | 0.962 | |
| RandomRotation | 0.793 | 0.832 | 0.813 | 0.847 | 0.804 | 0.898 | 0.840 | 0.903 | |
| ColorJitter | 0.778 | 0.836 | 0.805 | 0.874 | 0.814 | 0.907 | 0.854 | 0.927 | |
| RandomCrop+ RandomHorizontalFlip | 0.815 | 0.843 | 0.871 | 0.884 | 0.838 | 0.927 | 0.893 | 0.951 | |
| Avg | 0.796 | 0.848 | 0.842 | 0.885 | 0.816 | 0.916 | 0.859 | 0.933 | |
A.4 Experiments on Tiny-ImageNet Dataset
In Table 10, it can be observed that our method outperforms the random method in terms of attack success rates of poisoned samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively. In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases. When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method. These findings provide strong evidence supporting the effectiveness of our proposed method.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.952 | 0.990 | 0.983 | 0.989 | 0.960 | 0.992 | 0.985 | 0.991 | 0.956 | 0.991 | 0.981 | 0.989 |
| RandomCrop | 0.893 | 0.959 | 0.950 | 0.958 | 0.917 | 0.970 | 0.963 | 0.973 | 0.904 | 0.973 | 0.958 | 0.971 | |
| RandomHorizontalFlip | 0.952 | 0.989 | 0.984 | 0.989 | 0.963 | 0.991 | 0.981 | 0.989 | 0.947 | 0.983 | 0.974 | 0.985 | |
| RandomRotation | 0.881 | 0.932 | 0.937 | 0.944 | 0.912 | 0.959 | 0.964 | 0.966 | 0.910 | 0.957 | 0.963 | 0.970 | |
| ColorJitter | 0.956 | 0.990 | 0.985 | 0.990 | 0.966 | 0.993 | 0.986 | 0.992 | 0.956 | 0.990 | 0.983 | 0.988 | |
| RandomCrop+ RandomHorizontalFlip | 0.900 | 0.943 | 0.955 | 0.950 | 0.917 | 0.977 | 0.967 | 0.975 | 0.924 | 0.974 | 0.965 | 0.978 | |
| Avg | 0.922 | 0.967 | 0.966 | 0.970 | 0.939 | 0.980 | 0.974 | 0.981 | 0.933 | 0.978 | 0.971 | 0.980 | |
| Blended | None | 0.889 | 0.964 | 0.947 | 0.975 | 0.876 | 0.965 | 0.930 | 0.973 | 0.873 | 0.969 | 0.941 | 0.978 |
| RandomCrop | 0.836 | 0.928 | 0.920 | 0.942 | 0.835 | 0.931 | 0.914 | 0.936 | 0.822 | 0.927 | 0.906 | 0.932 | |
| RandomHorizontalFlip | 0.905 | 0.966 | 0.964 | 0.980 | 0.888 | 0.964 | 0.950 | 0.977 | 0.886 | 0.965 | 0.954 | 0.981 | |
| RandomRotation | 0.865 | 0.931 | 0.946 | 0.953 | 0.872 | 0.945 | 0.942 | 0.966 | 0.870 | 0.950 | 0.941 | 0.966 | |
| ColorJitter | 0.883 | 0.958 | 0.945 | 0.964 | 0.871 | 0.956 | 0.917 | 0.962 | 0.872 | 0.967 | 0.931 | 0.971 | |
| RandomCrop+ RandomHorizontalFlip | 0.850 | 0.931 | 0.940 | 0.958 | 0.840 | 0.929 | 0.919 | 0.942 | 0.828 | 0.931 | 0.912 | 0.940 | |
| Avg | 0.871 | 0.946 | 0.944 | 0.962 | 0.864 | 0.948 | 0.929 | 0.959 | 0.859 | 0.952 | 0.931 | 0.961 | |
| Optimized | None | 0.907 | 0.981 | 0.956 | 0.984 | 0.875 | 0.968 | 0.934 | 0.968 | 0.862 | 0.970 | 0.935 | 0.967 |
| RandomCrop | 0.897 | 0.976 | 0.959 | 0.980 | 0.886 | 0.968 | 0.950 | 0.972 | 0.869 | 0.958 | 0.945 | 0.951 | |
| RandomHorizontalFlip | 0.903 | 0.979 | 0.956 | 0.985 | 0.892 | 0.976 | 0.950 | 0.976 | 0.877 | 0.975 | 0.948 | 0.974 | |
| RandomRotation | 0.899 | 0.977 | 0.957 | 0.972 | 0.897 | 0.978 | 0.956 | 0.972 | 0.888 | 0.973 | 0.948 | 0.956 | |
| ColorJitter | 0.695 | 0.714 | 0.744 | 0.707 | 0.769 | 0.773 | 0.803 | 0.775 | 0.772 | 0.779 | 0.816 | 0.781 | |
| RandomCrop+ RandomHorizontalFlip | 0.892 | 0.973 | 0.960 | 0.974 | 0.890 | 0.964 | 0.952 | 0.968 | 0.870 | 0.963 | 0.949 | 0.957 | |
| Avg | 0.866 | 0.933 | 0.922 | 0.934 | 0.868 | 0.938 | 0.924 | 0.939 | 0.856 | 0.936 | 0.924 | 0.931 | |
A.5 Experiments on Different Pre-trained Feature Extractor
To compute the cosine similarity in the proposed PFS, a pre-trained feature extractor is required. This section examines the impact of various pre-trained feature extractors on the CIFAR-10 dataset. As demonstrated in Figure 6, even using a common feature extractor pre-trained on the ImageNet dataset yields significant improvements compared to the baseline.